OpenStudy (anonymous):

Andrea can clean a house in 4 times as fast as Denise. When they work together,Andrea & Denise can clean a large house in 8 hours. How many hours would it take Denise to clean the house by herself?

OpenStudy (anonymous):

@Loser66

hero (hero):

@Loujoelou, are you aware of equation: \(\dfrac{M \times N}{M + N} = t\) M = time it takes 1st person to do a job alone N = time it takes 2nd person to do a job alone t = time it takes both to complete the job together?

OpenStudy (anonymous):

no :) so if i used that equation would it be: 4x*x 4x^2 -----= ------ ? 4x+x 5x

hero (hero):

Yes, now what are you going to set that equal to?

OpenStudy (anonymous):

4x^2 4x ----- = t ---- = t ? 5x 5

OpenStudy (anonymous):

x= 5/4 ?

hero (hero):

How did you get that? Show the work you did to arrive at that result.

hero (hero):

You do realize a value for t is given, right?

OpenStudy (anonymous):

sure ok so i cancelled an x on the top & bottom of: (4x^2)/(5x) & omg i forgot about the 8 hours!

OpenStudy (anonymous):

(4x/5)= 8 so 8*(5/4)=10! So the answer is 10 right?

hero (hero):

Congrats.

OpenStudy (anonymous):

Thanks for guiding me through the problem :)

hero (hero):

Nice job solving.